OpenStudy (anonymous):

1,2,3,5,8,13,21 what is the pattern?

OpenStudy (anonymous):

You add the previous two. for example: 3=1+2 5=3+2 8=5+3 etc

OpenStudy (anonymous):

It is known as Fibonacci series

OpenStudy (anonymous):

Not quite. Fibonacci SEQUENCE has two 1's.